Determine the pH during the titration of 26.2 mL of 0.296 M perchloric acid by 0.200 M barium hydroxideat the following points:

(1) Before the addition of any barium hydroxide

(2) After the addition of 9.70 mL of barium hydroxide

(3) At the equivalence point

(4) After adding 25.2 mL of barium hydroxide
Answer #1

as given: volume of perchloric acid =26. 2mL = 26.28103 L Molarity of perch logic acid (MC104) = 0.296M Moles of 44104 = Mola22 volume of Bacon), = 9.70 ML = 9.70 x16 L Moles of Bacon? = Molarity of Bacon2x rolume of Bacon 0.200 mol x 9.70 xIBK = 0.0- pH = -log Cht] = -log(0.161983287) = 0.790529792 = 0.791(3) since Bacon is strong base and H. 104, a Strong add, pH at equivalence point = 7.00 (heutral) 643 volume of Bacoma = 25.2PH = -log[ut] = -log (0.0628 24 9027) = 1.27756129 = 1.28